OpenStudy (anonymous):

A ball is dropped from a height of 1.50m and rebounds to a height of 1.15m . Part A Approximately how many rebounds will the ball make before losing 93% of its energy? Express your answer as an integer.

OpenStudy (anonymous):

@ganeshie8 I got 4 and then i got 11 but there both wrong

ganeshie8 (ganeshie8):

try 10

OpenStudy (anonymous):

yess it was ten

OpenStudy (anonymous):

how did u get ten bc i got 11

ganeshie8 (ganeshie8):

how much % energy its losing when it hits the ground 1 time ?

ganeshie8 (ganeshie8):

(1.15/1.5)*100 = 23.33 %

ganeshie8 (ganeshie8):

so, we need to solve below thing :- (1-.233)^x = (1-.93)

OpenStudy (anonymous):

ohh I didnt do it that way

OpenStudy (anonymous):

how would you know how to solve this kind of problem?

ganeshie8 (ganeshie8):

how u did ?

ganeshie8 (ganeshie8):

may be it becomes clear to u, if u tell me how u did, and i spot the mistake... if any :)

OpenStudy (anonymous):

I did this 100-93=7% , then took 7% x 1.5= .105 this means that it has energy or something at 93% then 1.15/ 1/5 = .76666 so i just multiplied the 1st bounce then 2nd bounce and 3r bounce and kept on going until i got to 11 but i could be wrong bc it probably made a difference in sig figs

OpenStudy (anonymous):

I kept on multiplying by that number until I got to .105 and then counted how many bounces there were

OpenStudy (anonymous):

omg are we suppose to count the 1st drop at a bounce bc i did that and i dont think i was suppose to, ahhhh lol thank you so much

ganeshie8 (ganeshie8):

good logic :)

ganeshie8 (ganeshie8):

u shud have stopped at 10th bounce... thats all

OpenStudy (anonymous):

thank you so much =D